OpenStudy (anonymous):

A proportional relationship between the number of pounds of potatoes (x) and the price in dollars (y) is graphed, and the ordered pair (4, 3) is on the graphed line. Part A: What is the price of 1 pound of potatoes? Show your work, including the proportion you used to determine the price. (8 points) Part B: What does the ordered pair (8, 6) on the graph represent? Explain in words.

OpenStudy (anonymous):

okay, so we know the equation y=mx+b

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

but first

OpenStudy (anonymous):

we have to plug in (4,3) into that equation

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so that you can find b, right?

OpenStudy (anonymous):

and then, from there you can draw a line and find out how much it would cost for 1

OpenStudy (anonymous):

Ok can u stay to check my work?

OpenStudy (anonymous):

totally

OpenStudy (anonymous):

ill have some work too and ill snap a pic

OpenStudy (anonymous):

wait can you help me draw the graph?

OpenStudy (anonymous):

what do you mean? like the line?

OpenStudy (anonymous):

Yea like how would it be presented?

OpenStudy (anonymous):

okay so since we dont have a slope

OpenStudy (anonymous):

we know that slope is the change in y over the change in x

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

so we have 3=4m+b

OpenStudy (anonymous):

but we cant solve yet since we have two variables

OpenStudy (anonymous):

ok ok

OpenStudy (anonymous):

so we can do 3/4 as the slope using the given point right?

OpenStudy (anonymous):

so now we have

OpenStudy (anonymous):

\[3=\frac{ 3 }{ 4 }\times 4+b\]

OpenStudy (anonymous):

ok ok

OpenStudy (anonymous):

so what do we get when we solve that equation?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

or 0..

OpenStudy (anonymous):

well, 3/4 times 4 equals three

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

its zero

OpenStudy (anonymous):

so the y intercept is zero

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

so now

OpenStudy (anonymous):

we have the equation y=3/4x+0

OpenStudy (anonymous):

so we just plug in one for x

OpenStudy (anonymous):

and we get y is equal to 3/4

OpenStudy (anonymous):

ok

OpenStudy (anonymous):



OpenStudy (anonymous):

so now all i need to do is plug it i right?

OpenStudy (anonymous):

just graph for you to visualize

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

and then part b, means 8 pounds for 6 dollars

OpenStudy (anonymous):

ok thank you :) :) :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

your welcome! just tag if you need help again :)

OpenStudy (anonymous):

ok :)